You forecast the future cash flow of Jaffe Ltd, a young growing company. You expect fast growth through the first four years, at which time the company will mature and you expect it to grow by 3% per year forever more after. The following are the forecast cash flows:

Year 1: $5 million

Year 2: $10 million

Year 3: $15 million

Year 4: $20 million

Year 5: $20.6 million

Year 6: $20.218 million

Year 7 onward: continued growth at 3% per year

The discount rate is 9%. What is the value of Jaffe today at Year 0?

(Do not round intermediate calculations. Report your result in millions of dollars. Round the final answers to 2 decimal places. Omit $ sign and the word “million” in your response.)
Answer #1
Value of Jaffe today is the PV of the expected cash flows
when discounted at 9%:
Year Cash flow in $ million PVIF at 9% PV at 9%
1 $             5.00 0.91743 $                 4.59
2 $           10.00 0.84168 $                 8.42
3 $           15.00 0.77218 $             11.58
4 $           20.00 0.70843 $             14.17
$             38.76
Continuing value of cash flows = 20*1.03/(0.09-0.03) = $          343.33
PV of continuing value = 343.33*0.70843 = $         243.23
Value of Jaffe today $           281.98